OpenStudy (anonymous):

3 log2 (x-1) + log2 4=5

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

do you know that \(\log_2(4)=2\) since \(2^2=4\) ?

ganeshie8 (ganeshie8):

first work out \(\log_2 4\) : \[\large 4 = 2^?\]

OpenStudy (misty1212):

that means you have \[3\log_2(x-1)+2=5\\ 3\log(x-1)=3\] before you even start with those properties of the log

OpenStudy (misty1212):

then divide both sides by 3, and finally write in equivalent exponential form

OpenStudy (misty1212):

you got this from there ?

OpenStudy (anonymous):

The answer is 3

OpenStudy (anonymous):

In my book

OpenStudy (misty1212):

idk lets see \[3\log_2(x-1)=3\\ \log_2(x-1)=1\\ x-1=2^1\] yeah i guess \(x=3\)